OpenStudy (janu16):

Which logarithmic graph can be used to approximate the value of y in the equation 5^y = 12?

OpenStudy (janu16):

.

satellite73 (satellite73):

\[5^y=12\iff \log_5(12)=y\]

OpenStudy (janu16):

@satellite73 ik tht part but how d i graph it?

OpenStudy (janu16):

these are the options

OpenStudy (mathmate):

I assume you have a choice of different graphs. The graph you posted passes through the points (12,1), (1,0) which therefore is a graph of \(y=log_{12}(x)\). Recall that \(log_a(1)=0\) and \(log_a(a)=1\). Based on the above, choose the appropriate graph and post it for confirmation, based on \(y=log_5(12) \)

OpenStudy (janu16):

ya i posted all of them above

OpenStudy (mathmate):

What I meant is to post the answer that you choose.

OpenStudy (janu16):

when i solved y=log5(12) i got 1.543 so am i looking for that answer on the graph?

OpenStudy (janu16):

@mathmate

OpenStudy (mathmate):

That's one way to do it, and it will give you the correct choice. But the idea of the question is to choose the graph to estimate y in 5^y=12. Recall that \(log_a(1)=0\) and \(log_a(a)=1\). You would also get the same answer choice (graph) if you look for the graph which gives you \(log(5)=1\), in which case it will be the right graph to use. \(Then\) you use the graph to estimate y=\(log_5(12)\).

OpenStudy (janu16):

is it the first one?

OpenStudy (janu16):

@mathmate

OpenStudy (mathmate):

Reread what I just posted, and please explain why you chose the first one!

OpenStudy (janu16):

because the first one passes through 12,1 thats why i picked it

OpenStudy (mathmate):

If it passes through (12,1), what is the base of the log function?

OpenStudy (mathmate):

Recall that \(log_a(a)=1\)! Here you're looking at base=5.

OpenStudy (janu16):

yea the base is 5

OpenStudy (mathmate):

if log(12)=1, the log graph has base equal to 12!

OpenStudy (janu16):

yes so is it the last one ?

OpenStudy (janu16):

@mathmate

OpenStudy (mathmate):

Well, I have not seen much effort on your part, but I'll explain below and let you judge for yourself.  A typical log function (to base a) has a graph shown above, y=\(log_a(x)\) There is a vertical asymptote towards -inf when x=0, so log(0) is undefined. The graph is increasing for all values of x>0. For ANY base a >0, the graph of y=\(log_a(x)\) crosses the x-axis at x=1, i.e. \(log_a(1)=0\) for any a. Also, for ANY base a>0, the graph of y=\(log_a(x)\) crosses the line y=1 at x=a, i.e. \(log_a(a)=1\). The question is hinged upon the application of the last statement to locate the correct graph where a=5, which means that we are looking for the graph in which \(log_5(5)=1\). Now can you double check your answer with the last statement?

OpenStudy (mathmate):

 Look for the graph where a=5 and decide if you have the right choice.